A particle moves in a straight line such that its displacement, x meters, from a fixed point O on the line at time t seconds is given by 𝑥 = 40[𝑒 −2𝑡 − 𝑒 −4𝑡 ].


(a) Find the time when the particle is instantaneously at rest.


(b) Find the displacement of the particle from O when t = 3 s.


(c) Find the total distance travelled during the first 3 seconds of its motion


1
Expert's answer
2022-02-13T12:14:33-0500

Given:

"\ud835\udc65 (t)= 40[\ud835\udc52^{\u22122\ud835\udc61} \u2212 \ud835\udc52^{\u22124\ud835\udc61}]"


(a) Find the time when the particle is instantaneously at rest:

"x'(t)=40[-2\ud835\udc52^{\u22122\ud835\udc61}+4 \ud835\udc52^{\u22124\ud835\udc61}]=0"

"\ud835\udc52^{\u22122\ud835\udc61}(-2+4 \ud835\udc52^{\u22122\ud835\udc61})=0"

"\ud835\udc52^{\u22122\ud835\udc61}=1\/2"

"t=\frac{1}{2}\ln 2\:\rm s"

(b) Find the displacement of the particle from O when t = 3 s:

"d=x(3)=40[𝑒^{−6} − 𝑒^{−12}]=0.1\:\rm m"

(c) Find the total distance travelled during the first 3 seconds of its motion

"l=x(3)-x(0)=0.1-0=0.1\:\rm m"